OpenStudy (anonymous):

how do you solve x2 + 9x+20=0 by factoring?

OpenStudy (xguardians):

(x+4)(x+5)?

OpenStudy (anonymous):

(x+4)(x+5) = 0 Yea !

OpenStudy (xguardians):

Little trick I learned was that the outers of the parenthesis must equal the middle, and they must multiply to be the last number. In this case 4+5 = 9, and 4*5 = 20.